OpenStudy (anonymous):

write an equation in slope-intercept form of the line with parametric equations x=t-4, y= 3t+1

OpenStudy (anonymous):

Solve the x equation for t: t=x+4. Then plug it into y. y=3(x+4)+1 y=3x+12+1 y=3x+13